### v1

@@ 3.1 Data and features
Our data cover U.S. and international equity markets for the periods 1963 to 2018 and 1990 to 2018, respectively. We use the merged CRSP/Compustat database from the Wharton Research Data Service (WRDS) for the US sample and LSEG Datastream101010formerly called Reuters or Refinitiv Datastream. for the international sample. Data are pre-processed using the methodology described in Tobek and Hronec (2021). There are four regions in the sample - US, Europe, Japan and Asia-Pacific, which also includes 23 developed countries. [^11] US, Austria, Belgium, Denmark, Finland, France, Germany, Greece, Ireland, Italy, Luxembourg, Netherlands, Norway, Portugal, Spain, Sweden, Switzerland, UK, Japan, Australia, New Zealand, Hong Kong and Singapore. We report our results for both the full sample and the liquid sample, which includes only the most liquid stocks. The liquid sample includes stocks that meet the following criteria: a price greater than $1$ (or $0.1$ for Asia Pacific) at the end of the previous month, a market capitalisation within the top 95%, and a trading volume within the top 95%, and a trading volume within the top 95% of the total dollar trading volume over the previous 12 months in each region. The full sample includes stocks with a price greater than $1$ (\$0.1 for Asia Pacific) at the end of the previous month. Market capitalisation and volume filters are not applied to the full sample. Micro-cap stocks in the full sample therefore account for only a small fraction of the capitalisation and capitalisation and trading volume of the total market. The average, minimum and maximum number of stocks in the cross-section in each region is shown in table Table 1.
@@ 3.1.1 Features
The list of 153 anomalies is given in Appendix A and includes 93 fundamental, 49 market friction and 11 I/B/E/S anomalies, mostly covered by McLean and Pontiff (2016), Hou et al. (2020) and Harvey et al. (2016). Compared to the original data, we use weekly data during training. This means that The fundamental signals are updated every week with financial statement information from financial years ending at least 6 months earlier. Trade data information, such as market capitalisation, is taken as the most recent available. Missing characteristic values are imputed using cross-sectional medians.
Each sub-network in the two-stage quantile neural network model uses its own set of features. The standardized $\tau$-quantile sub-network uses 176 features consisting of 153 anomalies-based stock characteristics together with 18 stock-level volatilities and 5 market mean returns. The 18 stock-level volatilities are calculated using daily exponentially weighted moving averages (EWMA) with $\alpha\in\{0.8,0.9,0.94,0.96,0.98,0.99\}$ smoothing factors, EWMA of negative returns with $\alpha$ smoothing factors of $\{0.8,0.9,0.94\}$ to proxy asymmetric volatility, total volatilities calculated over the past 3, 6 and 12 months, and Parkinson (1980) high-low range-based volatility. All stock-level volatility estimates are scaled by their cross-sectional mean to remove the effects of changes in market-wide average volatility. This is done to make stock-level volatilities more consistent with the other stock-level variables, which are normalised to cross-sectional empirical quantiles. The 5 additional market variables used as inputs to the standardized $\tau$-quantile sub-network are computed as EWMAs of cross-sectional equal-weighted averages of individual stock returns $\mu_{t}=\alpha\cdot\left(1/N_{t}\sum_{i=1}^{N_{t}}r_{i,t-1}\right)+(1-\alpha)\cdot\mu_{t-1}$ over the number of stocks $N_{t}$ with smoothing parameters $\alpha\in\{0.9,0.94,0.96,0.99,0.999\}$. These variables measure the average historical market-wide return and should help in forecasting when there is some persistent time variance in expected market mean returns. We rescale the mean variables by the same scalar as the predicted returns to reflect the fact that their scale over time is relevant in the estimation.
The market volatility sub-network uses 18 cross-sectional averages of stock-level volatilities described above. These serve as a proxy for market-wide index volatility, as they are likely to be relevant for estimating the distribution of individual stock returns, which includes both systematic and idiosyncratic volatility components.
@@ 3.1.2 Validation
We use the initial period from 1973 to 1994 for hyperparameter optimization. Models with varying hyperparameters are trained from 1973 to 1989 and validated from 1990 to 1994. The hyperparameters yielding the best validation performance, measured in terms of average quantile loss (see Section 3.3 for details), are then applied to train models from 1995 to 2018, generating true out-of-sample predictions. Detailed hyperparameter searches for both full and liquid samples are documented in Appendix D together with the validation scheme.
From 1995 to 2018, using only the U.S. sample, we retrain the model annually using an expanding data window starting in 1973. To prevent data leakage into the next year, the latest date for feature calculation each year is November 29. Separate models are trained for full and liquid samples, with monthly out-of-sample predictions for both U.S. and international samples, using the most recently trained model at each month’s end. Predictions from the full sample model are applied to the full sample and, similarly, to the liquid sample.
@@ 3.2 Description of the estimated moments
Since our main empirical results are based on the predicted quantiles and moments of the distribution of all available stocks, we first describe the estimated moments for the US data. As the rest of the international regions show similar dynamics, we present these results in Appendix Table C1. Specifically, we predict 37 $\tau$ quantiles $\tau\in\{0.00005,0.0001,0.001,0.005$, $0.01,\ldots,0.05$, $0.075$, $0.1,0.15,0.2,\ldots,0.8,0.85,0.9$,$0.925$, $0.95,\ldots 0.99$,0.995,0.999,0.9999,0.99995}0.995,0.999,0.9999,0.99995\}0.995 , 0.999 , 0.9999 , 0.99995 }, mean, standard deviation, skewness and kurtosis of all available individual stock returns.
Table 2 contains summary statistics for the predicted moments, and Figure 4 shows the distribution of the predicted mean, standard deviation, skewness and kurtosis for the full and liquid samples for all stocks across the US.
Table 2: Summary statistics for the moments in the U.S. This table shows summary statistics for the forecasted first four moments for the full and liquid samples in the U.S. The moments are forecasted using the quantiles-to-moments algorithm introduced in Section 2.3.
@@
We can see that the distribution of the mean for both samples is approximately centred around zero, with the liquid sample having a slightly higher density around the mean. The mean return is positive for both samples, with the liquid sample having a slightly higher mean (0.0098) than the total sample (0.0068). The standard deviation is higher for the total sample (0.1477) than for the liquid sample (0.1071) and shows a wider spread for the total sample than for the liquid sample. Skewness is significantly higher in the full sample (0.8453) than in the liquid sample (0.1800), reflecting more extreme positive and negative returns, while the liquid sample shows a more symmetrical distribution. Kurtosis is also higher in the full sample (8.9784) than in the liquid sample (5.4261), indicating more pronounced tails compared to the liquid sample.
Figure 4: Histogram of Central Distribution Moments. This figure shows histograms of the forecasted first four moments for the full and liquid samples in the U.S. The Moments are forecasted using the quantiles-to-moments algorithm introduced in Section 2.3. Displayed variables are trimmed to exclude the top and bottom 0.5% of their values.
Table 3 also includes average cross-sectional correlations between the first four predicted moments and the median of stock returns. There is a high correlation of $0.95$ between the forecasted mean and the median for the liquid sample, while the correlation is lower ($0.75$) for the full sample. This is consistent with the fact that the full sample contains more micro-cap stocks with higher skewness, which affects the mean but not the median. There is also a high correlation between skewness and kurtosis ($0.91$) in the full sample. In contrast, in the liquid sample, the correlation between skewness and kurtosis is close to zero.
@@
Median 1.00 0.75 -0.63 -0.54 -0.44 1.00 0.95 -0.34 0.07 0.16
Mean 0.75 1.00 -0.22 -0.08 -0.08 0.95 1.00 -0.15 0.26 0.09
Variance -0.63 -0.22 1.00 0.53 0.24 -0.34 -0.15 1.00 0.66 -0.30
Skewness -0.54 -0.08 0.53 1.00 0.91 0.07 0.26 0.66 1.00 -0.05
Kurtosis -0.44 -0.08 0.24 0.91 1.00 0.16 0.09 -0.30 -0.05 1.00
@@ 3.3 Predictability
Here we evaluate the out-of-sample predictions of the quantiles from our two-stage quantile neural network. We also evaluate the predictions of the mean and variance estimated from the quantiles. This is of particular interest as we argue that the first and second moments, estimated without any assumptions about the distribution, are more predictive of the mean and variance. To benchmark our forecasting strategy, we use a linear model (LNN), the single hidden layer neural network with 32 neurons (1hNN), the double hidden layer neural network with 128 neurons (2hNN) and the commonly used GARCH. Details of the benchmark models can be found in Appendix B.1 and Appendix B.2.
To measure predictive performance across the distribution, we report the average loss, which is defined as the time series average of the cross-sectional $\tau$-quantile losses averaged across all $\tau$s for each month.
$L_{avg}=\frac{1}{T}\sum_{t=1}^{T}\left(\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\left(\frac{1}{K}\sum_{\tau\in\mathcal{T}}L_{\tau,t,i}\right)\right).$ (5)
@@
Predictive performance differences for the liquid sample mirror those observed for the full sample, although to a lesser extent and with fewer statistically significant differences. As expected, average out-of-sample quantile losses are generally lower for the liquid sample than for the full sample, reflecting the lower volatilities of the underlying stocks. The two-stage model consistently outperforms the GARCH model across all regions. However, when compared to alternative neural network specifications, the two-stage model only shows significant outperformance in the US region. Internationally, while the two-stage model achieves lower average quantile losses than the alternative neural network specifications, these differences do not reach statistical significance at the 5% level, except for the 1hNN model in Europe.
Table 4: Out-of-sample quantile forecasts evaluation. This table shows the average out-of-sample quantile cross-section loss for quantile forecasts from the two-stage model and alternative models. The results are obtained separately for the full and liquid samples for individual regions covering the out-of-sample period from 1995 to 2018. Neural network models are trained annually using the expanding window on US data only, while out-of-sample forecasts are generated for both the US and international samples. The average loss is calculated as the time series average of the cross-sectional averages of the average quantile loss across all $\tau\in\{0.00005,0.0001,0.001,0.005$, $0.01,\dots,0.05$, $0.075$, $0.1,0.15,0.2,\dots,0.8,0.85,0.9$,$0.925$, $0.95,\dots 0.99$,0.995,0.999,0.9999,0.99995}0.995,0.999,0.9999,0.99995\}0.995 , 0.999 , 0.9999 , 0.99995 } as in (Eq. 5) for the two-stage NN model, as well as benchmark neural network models (the linear model (LNN), the one-layer model (1hNN), the two-layer model (2hNN)) and the benchmark GARCH model. Note that losses are multiplied by 100.0, and the t-statistics in parentheses are computed for the average loss difference between the given model specification and the two-stage benchmark model using Newey-West standard errors with 12 lags.
@@
We also examine the performance of the mean and variance forecasts. To compare our two-stage model and the forecasts obtained by the algorithms described earlier, we compute the mean forecasts directly using a two-layer hidden neural network model estimated with a mean square error loss function. We follow Gu et al. (2020) and modify the calculation of the Diebold-Mariano test statistic to compare the cross-sectional average of forecast errors instead of comparing individual errors.
@@
Finally, to evaluate the second moment forecasts, we use the Mean Absolute Deviation (MAD) and the Root Mean Squared Error (RMSE) of the predicted 22-day volatility for the full and liquid samples for all regions. Table 6 shows a highly significant outperformance of the volatility forecast based on the two-stage model in terms of both MAD and RMSE for all regions and for both the full and liquid samples.
Table 6: Volatility forecast evaluation. This table shows the mean absolute deviation (MAD), root mean squared error (RMSE), and Diebold Mariano test statistic (DM) comparing the corresponding losses of the two competing models for the 22-day volatility forecasts. Volatility forecasting based on two-stage quantile forecasting (two-stage NN) is compared with GARCH model forecasting. Results are obtained separately for full and liquid samples on individual regions covering the period from 1995 to 2018. Neural network models are retrained annually using an expanding window on US data only, while out-of-sample forecasts are generated for both US and international samples. Diebold-Mariano test statistics are computed for the difference between the GARCH and two-stage NN model volatility forecasts. Note that losses are multiplied by 100.0 and the standard errors used in the Diebold-Mariano test statistics are adjusted using the Newey-West standard deviation estimator with 12 lags.
@@
Overall, we demonstrate the superior distributional forecasting performance of the two-stage model and the moments derived from it compared to alternative models. The average out-of-sample quantile loss represents the forecast over the entire distribution, while the evaluation of the mean and variance forecasts provides additional evidence derived from the forecasted quantiles.

### v2

@@ 3.1 Data and features
Our data cover U.S. and international equity markets for the periods 1963 to 2018 and 1990 to 2018, respectively. We use the merged CRSP/Compustat database from the Wharton Research Data Service (WRDS) for the US sample and LSEG Datastream888formerly called Reuters or Refinitiv Datastream. for the international sample. Data are pre-processed using the methodology described in Tobek and Hronec (2021). There are four regions in the sample - US, Europe, Japan and Asia-Pacific, which also includes 23 developed countries. [^9] US, Austria, Belgium, Denmark, Finland, France, Germany, Greece, Ireland, Italy, Luxembourg, Netherlands, Norway, Portugal, Spain, Sweden, Switzerland, UK, Japan, Australia, New Zealand, Hong Kong and Singapore. We report our results for both the full sample and the liquid sample, which includes only the most liquid stocks. The liquid sample includes stocks that meet the following criteria: a price greater than 111 (or 0.10.10.1 for Asia Pacific) at the end of the previous month, a market capitalisation within the top 95%, and a trading volume within the top 95%, and a trading volume within the top 95% of the total dollar trading volume over the previous 12 months in each region. The full sample includes stocks with a price greater than 111 (\$0.1 for Asia Pacific) at the end of the previous month. Market capitalisation and volume filters are not applied to the full sample. Micro-cap stocks in the full sample therefore account for only a small fraction of the capitalisation and capitalisation and trading volume of the total market. Table 1 summarizes the minimum, average, and maximum number of stocks per month observed in the cross-section of each region over the sample period, separately for the full and liquid samples.
@@ 3.1.1 Features
The list of 153 anomalies is given in Appendix A and includes 93 fundamental, 49 market friction and 11 I/B/E/S anomalies, mostly covered by McLean and Pontiff (2016), Hou et al. (2020) and Harvey et al. (2016). Compared to the original data, we use weekly data during training. This means that the fundamental signals are updated every week with financial statement information from financial years ending at least 6 months earlier. Trade data information, such as market capitalisation, is taken as the most recent available. Missing characteristic values are imputed using cross-sectional medians (Gu et al., 2020b).
[…]
The market-wide volatility sub-network uses 18 cross-sectional averages of stock-level volatilities described above. These serve as a proxy for market-wide index volatility, as they are likely to be relevant for estimating the distribution of individual stock returns, which includes both systematic and idiosyncratic volatility components.
@@ 3.1.2 Validation
We use the initial period from 1973 to 1994 for hyperparameter optimization. Models with varying hyperparameters are trained from 1973 to 1989 and validated from 1990 to 1994. The hyperparameters yielding the best validation performance, measured in terms of average quantile loss (see Section 3.3 for details), are then applied to train models from 1995 to 2018, generating out-of-sample predictions. Detailed hyperparameter searches for both full and liquid samples are documented in Appendix D together with the validation scheme.
From 1995 to 2018, using only the U.S. sample, we retrain the model annually using an expanding data window starting in 1973. To ensure that the 22-day ahead return [^10] Our forecast target., ends within the current calendar year, we restrict the final forecast date each year to November 29. All features are constructed using information available up to that point. This avoids using any information from the next year, thereby preventing both label and feature leakage from future data. Separate models are trained for full and liquid samples, with monthly out-of-sample predictions for both US and international samples, using the most recently trained model at each month’s end. Predictions from the full sample model are applied to the full sample and, similarly, to the liquid sample. Furthermore, to provide additional evidence of out-of-sample predictability robustness, we report predictability and profitability results for the period from 2019 to 2023 for the U.S. stocks in the Appendix F.
@@ 3.2 Description of the estimated moments
Since our main empirical results are based on the predicted quantiles and moments of the distribution of all available stocks, we first describe the estimated moments for the US data. As the rest of the international regions show similar dynamics, we present these results in Appendix Table C1. Specifically, we predict 37 $\tau$-quantiles with $\tau\in\mathcal{T}$, where $\mathcal{T}$ is the predefined set $\mathcal{T}=\{0.00005,0.0001,0.001,0.005,0.01,\ldots,0.05,0.075,0.1,0.15,0.2,\ldots,0.8,0.85,0.9,0.925,0.95,\\ \ldots,0.99,0.995,0.999,0.9999,0.99995\}$, which provides denser coverage in the tails and sparser spacing in the center of the distribution. This choice reflects a practical trade-off between statistical resolution and computational feasibility [^11] While using more quantiles is possible, it substantially increases computational and storage costs. We also experimented with fewer quantiles, but this led to inferior interpolation performance, particularly in the tails. Nonetheless, the architecture of the two-stage quantile neural network remains flexible and can accommodate changes in both the number and placement of quantiles—allowing, for example, denser tail coverage in applications such as Value-at-Risk.. The grid is deliberately designed to provide denser coverage in the tails of the distribution, where spline interpolation is more fragile and tail behaviour is critical for applications such as risk management while the central region is covered more sparsely (every 5%). This design also helps counteract the imbalance in loss aggregation since quantile losses are typically larger in the center, uniform averaging across $\tau$ implicitly downweights the tails. Denser tail coverage helps restore balance and ensures that tail performance is better accounted for.
Further, we also predict mean, standard deviation, skewness and kurtosis of all available individual stock returns. This pattern has practical relevance for portfolio construction. In the full sample, the high correlation implies that both characteristics yield similar asset ranking and thus similar products. In contrast, in the liquid sample, correlations close to zero support a potentially more diversified portfolio construction based on higher-order moments.
Table 2 contains summary statistics for the predicted moments, and Figure 4 shows the histogram of the predicted mean, standard deviation, skewness and kurtosis for the full and liquid samples for all stocks across the US.
Table 2: Summary statistics for the moments in the U.S. This table shows summary statistics for the forecasted first four moments for the full and liquid samples in the U.S. The moments are forecasted using the quantiles-to-moments algorithm introduced in Section 2.3. The table reports standard deviation, skewness, and kurtosis without bias adjustment described in Section E.3 as well as with the adjustment (Adj).
@@
Kurtosis Adj 13.798 9.9362 19.8491 6.8452 6.2727 8.3637
We can see that the distribution of the mean for both samples is approximately centred around zero, with the liquid sample having a slightly higher density around the mean. The mean return is positive for both samples, with the liquid sample having a slightly higher mean (0.0098) than the total sample (0.0068). The standard deviation is higher for the total sample (0.1477) than for the liquid sample (0.1071) and shows a wider spread for the total sample than for the liquid sample. Skewness is significantly higher in the full sample (0.9340) than in the liquid sample (0.2068), reflecting more extreme positive and negative returns, while the liquid sample shows a more symmetrical distribution. Kurtosis is also higher in the full sample (13.798) than in the liquid sample (6.8452), indicating more pronounced tails compared to the liquid sample. Bias adjustment described in Section E.3 has large impact only on Kurtosis, with negligible impact on standard deviation, and small impact on skewness.
Figure 4: Histogram of Central Distribution Moments. This figure shows histograms of the forecasted first four moments for the full and liquid samples in the U.S. The Moments are forecasted using the quantiles-to-moments algorithm introduced in Section 2.3 including bias adjustment described in Section E.3. Displayed variables are trimmed to exclude the top and bottom 0.5% of their values.
Table 3 also includes average cross-sectional correlations between the first four predicted moments and the median of stock returns. [^12] The table reports results with bias adjustment described in appendix Section E.3. The adjustment has only a small impact on the reported correlations. There is a high correlation of 0.950.950.95 between the forecasted mean and the median for the liquid sample, while the correlation is lower (0.750.750.75) for the full sample. This is consistent with the fact that the full sample contains more micro-cap stocks with higher skewness, which affects the mean but not the median. There is also a high correlation between skewness and kurtosis (0.910.910.91) in the full sample. In contrast, in the liquid sample, the correlation between skewness and kurtosis is close to zero. This pattern has practical relevance for portfolio construction. In the full sample, the high correlation implies that both characteristics yield similar asset ranking and thus similar products. In contrast, in the liquid sample, correlations close to zero support a potentially more diversified portfolio construction based on higher-order moments.
@@
Median 1.00 0.75 -0.63 -0.54 -0.43 1.00 0.95 -0.34 0.08 0.15
Mean 0.75 1.00 -0.22 -0.08 -0.08 0.95 1.00 -0.15 0.27 0.06
Var -0.63 -0.22 1.00 0.52 0.22 -0.34 -0.15 1.00 0.65 -0.39
Skewness -0.54 -0.08 0.52 1.00 0.91 0.08 0.27 0.65 1.00 -0.16
Kurtosis -0.43 -0.08 0.22 0.91 1.00 0.15 0.06 -0.39 -0.16 1.00
@@ 3.3 Predictability
This section evaluates the out-of-sample predictions of the quantiles from our two-stage quantile neural network. It also evaluates the predictions of the mean and variance estimated from the quantiles. This is of particular interest as we argue that the first and second moments, estimated without any assumptions about the distribution, are more predictive of the mean and variance. To benchmark our forecasting strategy, we use a linear model (LNN), the single hidden layer neural network with 32 neurons (1hNN), the double hidden layer neural network with 128 neurons (2hNN) and the commonly used GARCH. Details of the benchmark models can be found in Appendix B.1 and Appendix B.2. We also compare the performance of the two-stage quantile model with tree-based machine learning benchmarks in Appendix I. In addition, we provide a simulation study in Appendix H analyzing the performance of our forecasting strategy out of sample under a known data-generating process.
To measure predictive performance across the distribution, we report the average loss, which is defined as the time series average of the cross-sectional $\tau$-quantile losses, averaged across all $\tau$s for each month.
Specifically, we define the quantile loss for asset $i$ at time $t$ and quantile level $\tau$ as $L_{\tau,t,i}=\rho_{\tau}\left(r_{i,t+1}-\hat{Q}_{\tau}\right)$ as in Equation 4, where $r_{i,t+1}$ denotes the realized return, $\hat{Q}_{\tau}$ is the predicted $\tau$-quantile from the model.
We then compute the overall average quantile loss as
$L_{\text{avg}}=\frac{1}{T}\sum_{t=1}^{T}\left(\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\left(\frac{1}{K}\sum_{\tau\in\mathcal{T}}L_{\tau,t,i}\right)\right)$ (6)
where $T$ is the number of months, $N_{t}$ is the number of stocks available at time $t$, $K=|\mathcal{T}|$ is the number of quantiles considered, and $L_{\tau,t,i}$ is the quantile loss defined above.
@@
Predictive performance differences for the liquid sample mirror those observed for the full sample, although to a lesser extent and with fewer statistically significant differences. As expected, average out-of-sample quantile losses are generally lower for the liquid sample than for the full sample, reflecting the lower volatilities of the underlying stocks. The two-stage model consistently outperforms the GARCH model across all regions. However, when compared to alternative neural network specifications, the two-stage model only shows significant outperformance in the US region. Internationally, while the two-stage model achieves lower average quantile losses than the alternative neural network specifications, these differences do not reach statistical significance at the 5% level, except for the 1hNN model in Europe. In addition to aggregated quantile loss, we also report the $\tau$-specific out-of-sample quantile losses, in Table G1 and continuous ranked probability score in Table G2 in Appendix G.
These performance differences also highlight the role of model flexibility in capturing complex distributional patterns. In particular, the consistent improvement from LNN to 2hNN across regions reflects the benefits of modelling non-linearities in the conditional quantile functions. The two-stage model, which incorporates additional structural layers and nonlinear interactions across features and volatilities, achieves further gains—especially in the full sample—suggesting that the cross-sectional distribution of stock returns depends on non-linear effects that are not well captured by linear architectures.
Table 4: Out-of-sample quantile forecasts evaluation. This table shows the average out-of-sample quantile cross-section loss for quantile forecasts from the two-stage model and alternative models. The results are obtained separately for the full and liquid samples for individual regions covering the out-of-sample period from 1995 to 2018. Neural network models are trained annually using the expanding window on US data only, while out-of-sample forecasts are generated for both the US and international samples. The average loss is calculated as the time series average of the cross-sectional averages of the average quantile loss across all $\tau\in\mathcal{T}$, where $\mathcal{T}$ is defined in Section 3.2. The average loss is calculated as in (Eq. 6) for the two-stage NN model, as well as benchmark neural network models (the linear model (LNN), the one-layer model (1hNN), the two-layer model (2hNN)) and the benchmark GARCH model. Note that losses are multiplied by 100.0, and the t-statistics in parentheses are computed for the average loss difference between the given model specification and the two-stage benchmark model using Newey-West standard errors with 12 lags.
@@
We also examine the performance of the mean and variance forecasts. To compare our two-stage model and the forecasts obtained by the algorithms described earlier, we compute the mean forecasts directly using a two-layer hidden neural network model estimated with a mean square error loss function. We follow Gu et al. (2020a) and modify131313see Appendix G for the description of the testing procedure. the calculation of the Diebold-Mariano test statistic to compare the cross-sectional average of forecast errors instead of comparing individual errors.
@@
The results reflect a trade-off between robustness to extremes and sensitivity to the full shape of the conditional distribution. Quantile loss, by design, is less sensitive to large forecast errors, and the resulting median forecast is robust to outliers and heavy tails. However, this robustness can come at the cost of discarding potentially informative variation in the tails, which are particularly influential when forecasting the conditional mean—a non-robust functional. In the full sample, where extreme returns are more prevalent, this information appears valuable, and the mean forecast derived via numerical integration over the estimated quantile distribution outperforms both the median and the MSE-based benchmark. Notably, even the direct MSE-trained mean from the 2hNN achieves higher $R^{2}$ than the median in this setting, highlighting that robustness does not necessarily translate into superior predictive performance when the tails contain signal. In contrast, the liquid sample exhibits thinner tails and fewer extreme returns. In this setting, the benefits of robustness are more pronounced, and the median forecast achieves competitive or superior performance. Importantly, since both the mean and median are derived from the same underlying quantile forecasts in our two-stage model, the comparison isolates differences attributable to the target functional rather than the model class. These findings underscore that the relative performance of robust versus non-robust functionals depends on the tail behavior of the return distribution and the presence (or absence) of influential observations.
@@
Finally, to evaluate the second moment forecasts, we use the Mean Absolute Deviation (MAD) and the Root Mean Squared Error (RMSE) of the predicted 22-day volatility for the full and liquid samples for all regions. Realized volatility is calculated as the square root of the sum of squared daily returns over the 22 trading days. Table 6 shows a highly significant outperformance of the volatility forecast based on the two-stage model in terms of both MAD and RMSE for all regions and for both the full and liquid samples.
Table 6: Volatility forecast evaluation. This table shows the mean absolute deviation (MAD), root mean squared error (RMSE), and Diebold Mariano test statistic (DM) comparing the corresponding losses of the two competing models for the 22-day volatility forecasts. Volatility forecasting based on two-stage quantile forecasting (two-stage NN) is compared with GARCH model forecasting. Results are obtained separately for full and liquid samples on individual regions covering the period from 1995 to 2018. Neural network models are retrained annually using an expanding window on US data only, while out-of-sample forecasts are generated for both US and international samples. Diebold-Mariano test statistics are computed for the difference between the GARCH and two-stage NN model volatility forecasts. Note that losses are multiplied by 100 and the standard errors used in the Diebold-Mariano test statistics are adjusted using the Newey-West standard deviation estimator with 12 lags.
@@
To clarify the source of improvements, we emphasise that the accuracy of the empirical moment estimates—such as the mean and variance—primarily depends on the quality of the underlying distributional forecasts. While we rely on numerical integration over interpolated quantile forecasts, this approach is not tied to our two-stage neural network architecture. In principle, any sufficiently accurate quantile forecasting method, whether traditional or machine learning-based, can be used in conjunction with this integration technique to obtain distributional moments. However, our empirical results show that the two-stage QNN consistently delivers more accurate quantile forecasts than alternative models, particularly in the full sample. Overall, we demonstrate the superior distributional forecasting performance of the two-stage model and the moments derived from it compared to alternative models. The average out-of-sample quantile loss represents the forecast over the entire distribution, while the evaluation of the mean and variance forecasts provides additional evidence derived from the forecasted quantiles. While our evaluation focuses on distributional forecast accuracy and central moments, the underlying quantile estimates are also applicable to risk management tasks such as Value-at-Risk (VaR) forecasting. A formal backtesting exercise in this context is beyond the scope of this paper, but we acknowledge its practical relevance and view it as a promising direction for future work.